OpenStudy (anonymous):

Find the solution to the equation. e^5x=25

OpenStudy (anonymous):

is e an variable in your equation?

Nnesha (nnesha):

take ln both sides and (5x) is an exponent right ?

OpenStudy (anonymous):

yess

Nnesha (nnesha):

alright so you can take natural log both sides to cancel out e \[\huge\rm \cancel {ln e}^x = x \]

OpenStudy (anonymous):

ln25=5x?

Nnesha (nnesha):

\[\huge\rm \color{ReD}{ln }e ^{5x} = \color{Red}{\ln} 25\]

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

okay so now what?

Nnesha (nnesha):

use calculator for ln 25

Nnesha (nnesha):

and then solve for x

OpenStudy (anonymous):

i got 3.2188 when i did ln25

Nnesha (nnesha):

okay so 5x = 3.2188 solve for x divide both sides by 5

OpenStudy (anonymous):

0.6437?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

could you help me with one more?

Nnesha (nnesha):

i'll try

OpenStudy (anonymous):

2e^x-1=0

Nnesha (nnesha):

so exponent is (x-1) or just x ?

OpenStudy (anonymous):

just x

Nnesha (nnesha):

\[\huge\rm 2e^x -1 = 0\] first solve for e^x

OpenStudy (anonymous):

2e^x=1?

Nnesha (nnesha):

yes right now divide both sides by 2

OpenStudy (anonymous):

e^x=2?

Nnesha (nnesha):

no. what about one ?

Nnesha (nnesha):

2/1 = 2 but 1/2 isn't equal to 2

OpenStudy (anonymous):

0.5?

Nnesha (nnesha):

yes right right

Nnesha (nnesha):

e^x = .5 solve for x

OpenStudy (anonymous):

it would be .5, wouldnt it?

Nnesha (nnesha):

no how would you cancel out e ?

OpenStudy (anonymous):

ln.5=x

Nnesha (nnesha):

yes right ln.5 not just .5 :=)

OpenStudy (anonymous):

-0.6931

OpenStudy (anonymous):

okay this is the last one i promise! -3e^4x +11=2

Nnesha (nnesha):

right

Nnesha (nnesha):

same like the last one try it and let me know what you get i'll check it

OpenStudy (anonymous):

0.2747?

Nnesha (nnesha):

nice yes that's right :=)

OpenStudy (anonymous):

thank you! I get it now, but i just ran across this one that i dont know how to do it with the exponent like that , its e^x+1 =1.461

OpenStudy (anonymous):

i got 0.3791 so far but im stuck on dividing it by x+1

Nnesha (nnesha):

x is an exponent or x+1 is an exponent ?

OpenStudy (anonymous):

x+1

OpenStudy (anonymous):

is the answer -6.209?

Nnesha (nnesha):

alright you don't have to divide by x+1 remember to cancel out something thing from one side you have to apply opposite of addition/subtraction/division/multiplication so x+1 <--- here is x plus 1 adding so do opposite of one addition to cancel out one from left side

Nnesha (nnesha):

yes that's right! :-)